OpenStudy (anonymous):

What is the slope of the line passing through the points (2, 7) and (–1, 3)? And explain how to find the slope of the line

OpenStudy (anonymous):

A. 2/7 B. 3/4 C. 4/3 D. 1/3

OpenStudy (anonymous):

slope = rise/run = y2-y1/ x2-x1 = 3-7/-1-2 = -4/-3 = 4/3

OpenStudy (anonymous):

the answer should be C